हिंदी
तमिलनाडु बोर्ड ऑफ सेकेंडरी एज्युकेशनएचएससी विज्ञान कक्षा १२

Choose the correct alternative: If P(X = 0) = 1 – P(X = 1). If E[X] = 3 Var(X), then P(X = 0) is

Advertisements
Advertisements

प्रश्न

Choose the correct alternative:

If P(X = 0) = 1 – P(X = 1). If E[X] = 3 Var(X), then P(X = 0) is

विकल्प

  • `2/3`

  • `2/5`

  • `1/5`

  • `1/3`

MCQ
Advertisements

उत्तर

`1/3`

shaalaa.com
Mathematical Expectation
  क्या इस प्रश्न या उत्तर में कोई त्रुटि है?
अध्याय 11: Probability Distributions - Exercise 11.6 [पृष्ठ २२०]

APPEARS IN

सामाचीर कलवी Mathematics - Volume 1 and 2 [English] Class 12 TN Board
अध्याय 11 Probability Distributions
Exercise 11.6 | Q 11 | पृष्ठ २२०

संबंधित प्रश्न

For the random variable X with the given probability mass function as below, find the mean and variance.

`f(x) = {{:(1/10, x = 2","  5),(1/5, x = 0","  1","  3","  4):}`


For the random variable X with the given probability mass function as below, find the mean and variance.

`f(x) = {{:((4 - x)/6,  x = 1","  2","  3),(0,  "otherwise"):}`


For the random variable X with the given probability mass function as below, find the mean and variance.

`f(x) = {{:(2(x - 1), 1 < x ≤ 2),(0, "otherwise"):}`


For the random variable X with the given probability mass function as below, find the mean and variance.

`f(x) = {{:(1/2 "e"^(x/2),  "for"  x > 0),(0,  "otherwise"):}`


Choose the correct alternative:

Consider a game where the player tosses a six-sided fair die. If the face that comes up is 6, the player wins ₹ 36, otherwise he loses ₹ k2, where k is the face that comes up k = {1, 2, 3, 4, 5}. The expected amount to win at this game in ₹ is


Choose the correct alternative:

If X is a binomial random variable with I expected value 6 and variance 2.4, Then P(X = 5) is


Let X be a continuous random variable with probability density function
f(x) = `{{:(3/x^4",",  x ≥ 1),(0",",  "otherwise"):}`
Find the mean and variance of X


What are the properties of Mathematical expectation?


State the definition of Mathematical expectation using continuous random variable


In a business venture a man can make a profit of ₹ 2,000 with a probability of 0.4 or have a loss of ₹ 1,000 with a probability of 0.6. What is his expected, variance and standard deviation of profit?


The number of miles an automobile tire lasts before it reaches a critical point in tread wear can be represented by a p.d.f.
f(x) = `{{:(1/30 "e"^(- x/30)",",  "for"  x > 0),(0",",  "for"  x ≤ 0):}`
Find the expected number of miles (in thousands) a tire would last until it reaches the critical tread wear point


A person tosses a coin and is to receive ₹ 4 for a head and is to pay ₹ 2 for a tail. Find the expectation and variance of his gains


Choose the correct alternative:

E[X – E(X)] is equal to


Choose the correct alternative:

E[X – E(X)]2 is


Choose the correct alternative: 

`int_(-oo)^oo` f(x) dx is always equal to


Choose the correct alternative: 

An expected value of a random variable is equal to it’s


Choose the correct alternative: 

A discrete probability function p(x) is always non-negative and always lies between


What is the expected value of a game that works as follows: I flip a coin and if tails pay you ₹ 2; if heads pay you ₹ 1. In either case, I also pay you ₹ 0.50


Share
Notifications

Englishहिंदीमराठी


      Forgot password?
Use app×